# Improving Image Captioning Descriptiveness by Ranking and LLM-based Fusion

### v1

@@ 4 Analysis
Our analysis to verify the effectiveness of the proposed model is performed on the widely used MS-COCO Karpathy split Karpathy and Fei-Fei (2015), which includes 5000 images.
@@ 4.1 Quantitative analysis
Quality evaluation.
In order to assess the quality of the generated captions, we employ four widely used image-agnostic evaluation metrics, i.e., BLEU@4 (B@4), METEOR (M), CIDEr (C), and SPICE (S). Additionally, we exploit the defined BLIPScore to further evaluate the performance of the generated captions. The BLIPScore allows to estimate the semantic alignment between the generated captions and the provided image.
Diversity evaluation.
@@
To investigate the diversity of the generated captions, we use mBLEU, and n-gram diversity (i.e., Div-$n$ Aneja et al. (2019)). These metrics evaluate diversity by comparing the $n$-gram differences among the generated captions that belong to the same image.
Richness evaluation.
@@
Part-of-Speech (POS) plays a crucial role in analyzing text semantics. To compare the semantic richness of the captions generated by the different models, we adopted POS tagging. In this regard, the Universal tagset consisting of 17 English POS tags is exploited Smith and Eisner (2005). Initially, the words in each caption are annotated with their respective POS tags. Subsequently, we estimate the frequency of each POS tag within the captions generated by each model.
Results.
@@
In our analysis we include the five literature models considered for caption generation, namely BLIP-2, ExpNet-v2, GIT, OFA, and ViT-GPT2. The performance of these models is compared with that of our proposed model, denoted as “Our (Fusion)”. Furthermore, we evaluate the performance achieved by selecting the caption with the highest BLIPScore from the five generated captions during the ranking phase for each image. The results obtained using this approach are presented as “Our (Best)”.
Several observations can be drawn from the quality and diversity results presented in Table 2. First, the performance of our proposed model in terms of captioning quality metrics is either on par with or slightly lower than existing models. It is worth noting that there is a noticeable decline in performance compared to the best-performing method, which is OFA, specifically for metrics B@4 and C, where our model lags behind with scores of 14 and 50, respectively. However, our model achieves comparable results to OFA for metrics M and S. Second, the proposed model is able to generate captions that describe the image more adequately. The BLIPScore is in fact 0.05 higher than the first SoTA model, which is OFA. We also highlight that the “Our (Best)” approach also obtains a BLIPScore equal to that of “Our (Fusion)”. In Appendix B.1, we provide a detailed analysis of the BLIPScore, including statistics on the models to which the selected captions for the fusion step belong. Third, diversity metrics demonstrate how the proposed model generates captions that are very different from the others. This difference is particularly noticeable in the case of Div-1 and Div-2, where the gap with the leading SoTA captioner is 0.15. Regarding mBLEU, our proposed model ranks second to ViT-GPT2. However, it is worth noting that ViT-GPT2 achieves higher diversity at the expense of faithfully and fluently describing the content, as evidenced by its significantly lower quality captioning performance.
@@
Table 2: Captioning performance in terms of image-agnostic metrics together with diversity statistics on the Karpathy split of MS-COCO. For each metric, best and second best models are highlighted in bold and underlined, respectively.
Model Quality Diversity
B@4 ($\uparrow$) M ($\uparrow$) C ($\uparrow$) S ($\uparrow$) BLIPScore ($\uparrow$) mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$)
BLIP-2 43.8 31.7 145.9 25.2 0.64 0.88 0.50 0.32
ExpNet-v2 40.7 30.0 139.5 24.4 0.59 0.84 0.49 0.32
GIT 38.7 29.5 131.2 23.3 0.62 0.80 0.49 0.32
OFA 44.6 32.5 153.7 26.6 0.65 0.91 0.51 0.33
ViT-GPT2 35.4 27.9 119.1 21.2 0.48 0.69 0.47 0.30
Our (Best) 41.6 31.6 145.5 25.9 0.70 1.00 0.52 0.34
Our (Fusion) 30.0 31.5 106.0 26.5 0.70 0.72 0.66 0.44
Table 3 shows the results of the comparison between the captioning models in terms of average number of tokens and frequency of POS-Tags. The first consideration that can be made is that the captions generated by our model are on average longer both than those generated by SoTA models and those composed by humans. Although this finding alone may not hold significant meaning, a closer examination of POS-Tag frequencies, particularly adjectives, nouns, and verbs, reveals that our captions exhibit a greater richness and descriptive quality.
@@
Table 3: Number of tokens and frequency of POS tags for generated captions. Each column reports each captioner’s statistics for the 5000 images of the MS-COCO test set (tags that never occurred are omitted). The “Ground-Truth” column indicates statistics for human generated captions. For each tag, best and second best models are highlighted in bold and underlined, respectively.
@@
Token 9.62$\pm$1.42 9.57$\pm$1.75 9.60$\pm$1.59 9.81$\pm$1.28 9.17$\pm$1.53 10.05$\pm$1.57 13.10$\pm$3.32 10.43$\pm$2.36
Adjective 0.62$\pm$0.81 0.52$\pm$1.12 0.51$\pm$0.74 0.55$\pm$0.75 0.39$\pm$0.68 0.66$\pm$0.81 0.91$\pm$0.98 0.85$\pm$0.91
Adposition 1.67$\pm$0.85 1.74$\pm$0.83 1.67$\pm$0.85 1.71$\pm$0.84 1.54$\pm$0.79 1.74$\pm$0.86 2.10$\pm$1.11 1.68$\pm$0.95
Adverb 0.01$\pm$0.10 0.00$\pm$0.03 0.01$\pm$0.10 0.00$\pm$0.07 0.01$\pm$0.07 0.01$\pm$0.09 0.06$\pm$0.26 0.07$\pm$0.28
Conjunction 0.20$\pm$0.41 0.21$\pm$0.44 0.23$\pm$0.44 0.24$\pm$0.45 0.23$\pm$0.43 0.27$\pm$0.47 0.46$\pm$0.59 0.24$\pm$0.47
Determiner 2.37$\pm$0.70 2.50$\pm$0.70 2.41$\pm$0.72 2.45$\pm$0.76 2.40$\pm$0.72 2.44$\pm$0.77 2.93$\pm$1.02 2.21$\pm$0.95
Noun 3.51$\pm$0.85 3.59$\pm$0.84 3.52$\pm$0.90 3.64$\pm$0.81 3.43$\pm$0.82 3.71$\pm$0.89 4.59$\pm$1.42 3.69$\pm$1.15
Numeral 0.09$\pm$0.29 0.11$\pm$0.32 0.09$\pm$0.29 0.14$\pm$0.35 0.05$\pm$0.22 0.14$\pm$0.35 0.16$\pm$0.42 0.12$\pm$0.35
Particles 0.18$\pm$0.40 0.13$\pm$0.35 0.15$\pm$0.39 0.15$\pm$0.36 0.13$\pm$0.36 0.16$\pm$0.39 0.19$\pm$0.42 0.19$\pm$0.43
Pronouns 0.09$\pm$0.29 0.06$\pm$0.24 0.12$\pm$0.35 0.09$\pm$0.28 0.09$\pm$0.29 0.09$\pm$0.31 0.14$\pm$0.37 0.14$\pm$0.38
Verb 0.85$\pm$0.61 0.71$\pm$0.57 0.89$\pm$0.72 0.85$\pm$0.67 0.90$\pm$0.75 0.84$\pm$0.68 1.55$\pm$0.95 1.21$\pm$0.89
@@ 4.2 Subjective study
We conduct a subjective study to collect human judgments for selecting the best caption. Specifically, participants were instructed to choose the caption that accurately describes the image, is grammatically correct, contains no incorrect information, is relevant to the image, and is human-like. Participants could select from six different captions generated by using BLIP-2, ExpNet-v2, GIT, OFA, ViT-GPT2, and the proposed model. To avoid bias, captions are randomly arranged each time the page is loaded. We use the Amazon Mechanical Turk (AMT) to gather the human judgements. From the MS-COCO test set, a subset of 1000 images is randomly sampled for evaluation. We conduct two distinct studies: the first involves rating each image by three unique workers, who are not necessarily domain experts; in the second study, we enlist the expertise of three domain experts to rate the images. Figure 4(a) displays a screenshot of the AMT interface designed for caption assessment. The interface is self-contained as instructions to guide the selection are given in the evaluation of each image. Table 4(b) shows the percentage of votes obtained by each model from generic workers (shown in the “Worker votes” column), expert workers (reported in the column “Expert votes”), and the average percentage between the two categories of participants (see column “Average votes”). Our captions have been selected more frequently by both categories of participants with a percentage of 37.77. The second best is OFA with a percentage of 16.18. Non-expert workers show a uniform distribution in the selection of SoTA captions (on average about 15%), while instead they show a greater, albeit limited, preference for our captions with a percentage of 24.33. It is worth noting that experts have consistently chosen our captions, with 51.20% of the votes against an average of about 10%.
Model Worker votes (%) Expert votes (%) Average votes (%)

### v3

@@ 4 Analysis
To assess the effectiveness of the proposed model, we conduct experiments on the widely used MS-COCO dataset, which consists of 82,783 training images and 40,504 validation images, each annotated with five descriptive captions [18]. We also evaluate our model on the Flickr30k dataset, an extended version of Flickr8k, containing 31,783 images, each paired with five captions [49]. To ensure consistency with state-of-the-art methods, we adopt the widely used data splits introduced by Karpathy and Li [18], utilizing 5,000 test images for MS-COCO and 1,000 test images for Flickr30k.
4.1 Qualitative results
Figure 6 showcases images accompanied by their respective ground-truth and generated captions. From the provided examples, it is evident that our model excels in producing more detailed captions. For instance, in the image featuring a child with a toothbrush, our model seamlessly integrates information about the toothbrush, overalls, and hands. Similarly, in the image of the tower with a clock, our model enriches the description with intriguing qualitative details such as the presence of red and gold. However, consistent with the discussions presented in the section outlining the limitations of our model, when the captions generated by the SoTA models exhibit significant similarity, our generated captions may not provide additional details (as observed in the image featuring the pole on the sidewalk). In some cases, the generated caption may also include redundant information, as seen in the description of the ducks in the pond, where terms like “muddy pond” and “brown water” are used together.
Ground-truth captions: A street sign for queen street with a colorful representation of the street on the pole. A decorated street sign for Queen Street with two blue signs below. A close up a street pole with a homemade street sign. We are looking at a street sign at the corner. A pole holding the street sign for Queen Street is decorated with a painting of a queen. Captions generated by existing models: BLIP-2: A street sign on a pole on the side of the road. ExpNet-v2: A wooden pole with street signs on the side of it. GIT: A street sign on a pole with a blue sign on it. OFA: A street sign on a pole on the side of a street. ViT-GPT2: A street sign on a pole on a sidewalk. Caption generated by the proposed model: Our: A wooden pole with street signs mounted on the sidewalk.
Ground-truth captions: A infant holding a baby toothbrush in his hand looking at it. A small baby is holding a white and blue toothbrush. A child holds a toothbrush in their hand. The baby is holding and looking at his tooth brush. A baby holds a toothbrush in its hand. Captions generated by existing models: BLIP-2: A baby in overalls holding a toothbrush. ExpNet-v2: A baby holding a toothbrush in his hand. GIT: A baby holding a toothbrush in his hands. OFA: A baby holding a toothbrush in its hands. ViT-GPT2: A baby holding a toothbrush in its mouth. Caption generated by the proposed model: Our: A baby in overalls is holding a toothbrush in his hands.
Ground-truth captions: Two Mallards float and drink water in a shallow pool. Two ducks in a pond in a camp grounds. Two ducks swimming in a body of water. Two ducks swimming on a very dirty pond. Two ducks are swimming in some brown water. Captions generated by existing models: BLIP-2: Two ducks swimming in a pool of brown water. ExpNet-v2: Two ducks swimming in the water in a pond. GIT: Two ducks swimming in a muddy pond with a car in the background. OFA: A couple of ducks swimming in a body of water. ViT-GPT2: A duck swimming in a pond with a duck in it. Caption generated by the proposed model: Our: Two ducks swimming in a muddy pond with a car in the background, amidst a pool of brown water.
Ground-truth captions: A large tall tower with a clock on the top. A red, white and gold decorative pillar with a clock on top. Pedestrians rest at the base of a large, public ornamental clock. A large tower in a crowded area with a clock at the top. A tall red and white clock tower in the middle of a street. Captions generated by existing models: BLIP-2: A red and gold tower with a clock on top. ExpNet-v2: A red clock tower in the middle of a street. GIT: A large clock tower with a large clock on it. OFA: A red clock tower with people sitting around it. ViT-GPT2: A clock tower with a clock on it’s side. Caption generated by the proposed model: Our: A red and gold clock tower with people sitting around it.
@@ 4.2 Quantitative analysis
In this section, we quantitatively assess the effectiveness of the proposed model. Specifically, we evaluate the quality of the generated captions and compare the results against existing models. We also analyze the diversity and richness of the captions, particularly in relation to those produced by selected captioning models.
4.2.1 Quality evaluation
To assess the quality of the generated captions, we utilize four widely adopted image-agnostic metrics: BLEU@4 (B@4), METEOR (M), CIDEr (C), and SPICE (S). These traditional metrics compare the tested captions to reference captions – assuming the latter as ideal descriptions – they may fall short in scenarios that emphasize descriptiveness. This limitation arises because existing datasets often contain reference captions that lack comprehensive details about images. As a result, $n$-gram-based metrics may not effectively measure or promote enhanced descriptiveness [54]. To address this issue, we incorporate ALOHa [33], CAPTURE [11], and Polos [43] to assess caption-image alignment and hallucination reduction. ALOHa is an open-vocabulary evaluation metric designed to detect object hallucinations in image captions by leveraging LLMs. Unlike traditional metrics such as CHAIR [35], which are constrained to a predefined list of objects and synonyms, ALOHa operates without such limitations. It extracts potentially groundable objects from candidate captions using an LLM, computes their semantic similarity to reference objects—sourced from human-written captions and image-based object detections—and applies Hungarian matching to align entities. The final hallucination score reflects the proportion of mismatched objects, enabling a more flexible and comprehensive assessment of object-level fidelity. CAPTURE is an automatic evaluation metric that assesses image descriptions by identifying key visual elements. It employs a scene graph parser to extract objects, attributes, and relationships from both candidate and reference captions. To enhance robustness, abstract nouns are filtered using a stop-word list. Evaluation is based on the F1 score, computed using exact matches, synonym matches, and soft semantic matches, thereby aligning well with human judgment. The final CAPTURE score is a weighted combination of the F1 scores for objects, attributes, and relationships:
$\text{CAPTURE}=\frac{\alpha\cdot F1_{\text{obj}}+\beta\cdot F1_{\text{attr}}+\gamma\cdot F1_{\text{rel}}}{\alpha+\beta+\gamma},$ (1)
where $\alpha=5$, $\beta=5$, and $\gamma=2$. Polos is a supervised evaluation metric that processes multimodal inputs through a parallel feature extraction architecture. It utilizes embeddings derived from large-scale contrastive learning to better capture semantic alignment between images and captions. Polos is specifically designed to correlate strongly with human evaluations and handle a wide variety of visual and linguistic content.
4.2.2 Diversity evaluation
@@
To investigate the diversity of the generated captions, we use mBLEU, and n-gram diversity (i.e., Div-$n$ [4]). These metrics evaluate diversity by comparing the $n$-gram differences among the captions generated that belong to the same image.
4.2.3 Richness evaluation
@@
Part-of-Speech (POS) plays a crucial role in analyzing text semantics. To compare the semantic richness of the captions generated by the different models, we adopted POS tagging. In this regard, the Universal tagset consisting of 17 English POS tags is exploited [40]. Initially, the words in each caption are annotated with their respective POS tags. Subsequently, we estimate the frequency of each POS tag within the captions generated by each model.
4.2.4 Results
@@
In our analysis, we include several literature models for caption generation, namely BLIP-2, EVCap, ExpNet-v2, GIT, I-Tuning, MSRM, OFA, VIPCap, and ViT-GPT2. The performance of these models is compared to that of our proposed fusion-based model, denoted as “Our (Fusion)”. In contrast, “Our (Best)” refers to the selection-based strategy in which, for each image, we retain the single caption among the five candidates that receives the highest BLIPScore during the ranking phase.
@@
Table 2: Captioning performance in terms of image-agnostic metrics on the Karpathy split of MS-COCO. For each metric, best and second best models are highlighted in bold and underlined, respectively. “–” indicates that the model has not reported a score on the metric.
Model B@4 M C S CAPTURE $F1_{\text{attr}}$ $F1_{\text{obj}}$ $F1_{\text{rel}}$
BLIP-2 43.8 31.7 145.9 25.2 0.36 10.96 61.02 38.85
EVCap 41.5 31.2 140.1 24.7
ExpNet-v2 40.7 30.0 139.5 24.4 0.36 9.47 61.26 40.75
GIT 38.7 29.5 131.2 23.3 0.39 10.75 64.15 46.98
I-Tuning 34.8 28.3 116.7 21.8
MSRM 41.7 30.7 134.8
OFA 44.6 32.5 153.7 26.6 0.38 11.34 62.99 42.88
ViPCap 37.7 28.6 122.9 21.9
ViT-GPT2 35.4 27.9 119.1 21.2 0.35 6.50 61.43 40.55
Our (Best) 41.6 31.6 145.5 25.9 0.39 12.93 62.70 42.68
Our (Fusion) 30.0 31.5 106.0 26.5 0.42 16.75 64.59 47.77
Table 3: Captioning performance in terms of image-agnostic metrics on the Karpathy split of Flickr30k. For each metric, best and second best models are highlighted in bold and underlined, respectively. “–” indicates that the model has not reported a score on the metric.
Model B@4 M C S CAPTURE $F1_{\text{attr}}$ $F1_{\text{obj}}$ $F1_{\text{rel}}$
BLIP-2 31.9 25.2 85.9 18.5 0.33 10.29 55.93 34.12
EVCap 84.4 18.0
ExpNet-v2 24.5 21.1 60.9 15.2 0.34 9.48 57.63 37.21
GIT 28.0 23.4 76.3 16.9 0.36 11.65 59.37 40.30
MSRM 28.9 25.4 58.3
I-Tuning 72.3 19.0
OFA 27.6 22.9 74.4 16.6 0.32 8.17 54.53 33.17
ViPCap 66.8 17.2
ViT-GPT2 16.4 17.9 39.2 11.1 0.30 7.91 52.14 28.80
Our (Best) 31.3 25.5 88.3 19.2 0.36 13.10 58.29 38.01
Our (Fusion) 24.0 26.5 73.1 21.0 0.40 18.29 60.70 43.11
Several observations can be drawn from the results presented in Tables 2, 3, 4, and 5. While the proposed model does not consistently outperform SoTA captioners in traditional quality metrics such as BLEU-4 (B@4) and CIDEr (C), it shows competitive or superior performance in metrics that better capture semantic alignment and linguistic richness. Specifically, “Our (Fusion)” achieves the highest SPICE score among all models on both datasets (26.5 on MS-COCO and 21.0 on Flickr30k), indicating strong alignment with the image content in terms of scene elements and relationships. Although OFA and BLIP-2 lead in B@4 and CIDEr on both benchmarks, our model obtains top results in SPICE and METEOR, suggesting that the captions it generates are more semantically faithful and fluent, even if they differ in surface n-gram similarity from the references. “Our (Best)” also performs strongly on traditional metrics—achieving the best CIDEr score (88.3) and second-best scores in B@4 and SPICE on Flickr30k—highlighting the effectiveness of candidate selection when constrained to single-caption outputs. Importantly, the fusion-based strategy leads to a substantial improvement in caption diversity. As shown in Table 4, “Our (Fusion)” achieves the highest Div-1 and Div-2 scores on both datasets (0.66/0.44 on MS-COCO and 0.61/0.40 on Flickr30k), significantly outperforming all baselines. It also achieves the second-lowest mBLEU, indicating low redundancy across generated captions while preserving quality. In contrast, models such as ViT-GPT2 exhibit similar mBLEU scores but suffer from substantially lower content quality, underscoring that high diversity alone does not guarantee caption adequacy. Moreover, the linguistic analysis in Table 5 reveals that our model generates longer and more syntactically complex sentences, with the highest frequencies of adjectives, nouns, verbs, and conjunctions. This indicates an increase in both lexical richness and descriptive capacity, beyond what is captured by standard metrics. Interestingly, the fusion model even surpasses human-written references in several of these linguistic dimensions, suggesting that it does not simply interpolate across candidates but produces more nuanced and detailed descriptions. Finally, results in Table 6 confirm that these advantages generalize across datasets and metrics, with “Our (Fusion)” achieving the best performance on both ALOHa and Polos scores for MS-COCO and Flickr30k. These consistent gains across quality, diversity, and semantic grounding metrics demonstrate the effectiveness of the proposed fusion-based captioning strategy in addressing limitations of existing models, particularly those reliant on single-decoder architectures.
Table 4: Captioning performance in terms of diversity statistics on the Karpathy split of MS-COCO and Flickr30k. For each metric, best and second best models are highlighted in bold and underlined, respectively.
Model MS-COCO Flickr30k
mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$) mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$)
BLIP-2 0.88 0.50 0.32 0.68 0.44 0.28
ExpNet-v2 0.84 0.49 0.32 0.52 0.41 0.26
GIT 0.80 0.49 0.32 0.60 0.42 0.27
OFA 0.91 0.51 0.33 0.64 0.43 0.27
ViT-GPT2 0.69 0.47 0.30 0.35 0.41 0.27
Our (Best) 1.00 0.52 0.34 1.00 0.54 0.35
Our (Fusion) 0.72 0.66 0.44 0.50 0.61 0.40
@@
Table 5: Number of tokens and frequency of POS tags for generated captions. Each column reports each captioner’s statistics for the 5000 images of the MS-COCO test set (tags that never occurred are omitted). The “Ground-Truth” column indicates statistics for human-generated captions. For each tag, best and second best models are highlighted in bold and underlined, respectively.
@@
[…]
Table 6: ALOHa and Polos results on the considered datasets. For each metric, best and second best models are highlighted in bold and underlined, respectively.
Model MS-COCO Flickr30k
ALOHa Polos ALOHa Polos
BLIP-2 0.73 0.71 0.61 0.66
ExpNet-v2 0.71 0.69 0.56 0.58
GIT 0.70 0.70 0.61 0.65
OFA 0.73 0.72 0.57 0.63
ViT-GPT2 0.68 0.66 0.47 0.47
Our (Best) 0.69 0.72 0.59 0.67
Our (Fusion) 0.75 0.74 0.60 0.69
@@ 4.3 Subjective study
We conduct a subjective study to collect human judgments for selecting the best caption. Specifically, participants were instructed to choose the caption that accurately describes the image, is grammatically correct, contains no incorrect information, is relevant to the image, and is human-like. Participants could select from six different captions generated by using BLIP-2, ExpNet-v2, GIT, OFA, ViT-GPT2, and the proposed model. To avoid bias, captions are randomly arranged each time the page is loaded. We use the Amazon Mechanical Turk (AMT) to gather the human judgments. From the MS-COCO test set, a subset of 1000 images is randomly sampled for evaluation. We conduct two distinct studies: the first involves rating each image by three unique workers, who are not necessarily domain experts; in the second study, we enlist the expertise of three domain experts to rate the images. Figure 8(a) displays a screenshot of the AMT interface designed for caption assessment. The interface is self-contained, as instructions to guide the selection are given in the evaluation of each image. Table 8(b) shows the percentage of votes obtained by each model from generic workers (shown in the “Worker votes” column), expert workers (reported in the column “Expert votes”), and the average percentage between the two categories of participants (see column “Average votes”). Our captions have been selected more frequently by both categories of participants, with a percentage of 37.77. The second best is OFA with a percentage of 16.18. Non-expert workers show a uniform distribution in the selection of SoTA captions (on average about 15%), while instead they show a greater, albeit limited, preference for our captions with a percentage of 24.33. It is worth noting that experts have consistently chosen our captions, with 51.20% of the votes against an average of about 10%.
Figure 7 illustrates the level of agreement among the six workers involved in the subjective study when selecting the best caption for each image. The graph addresses the question of how many of the 1000 images received a unanimous consensus among the workers. As observed from the results, there is a general consensus among the workers that our captions are superior. Specifically, for 308 images, two workers agreed on our caption as the best, for 266 images, three workers agreed, for 118 images, four workers agreed, and finally, for 31 images, five workers chose our caption.
Model
Worker
votes (%)
Expert
votes (%)
Average
votes (%)
@@
4.4 Ablation study
In this section, we present an ablation study on key components of the proposed model. Specifically, we analyze: (i) caption selection, focusing on BLIPScore variations among generated captions and the optimal number of captions to select; (ii) the impact of using alternative LLMs for caption fusion; and (iii) the effect of different prompt engineering strategies for caption fusion.
4.4.1 Caption selection
In this section, we first examine the rationale behind selecting only the top two captions rather than multiple captions. We then provide details on the methods used to generate the most frequently selected captions.
Figure 9 illustrates the distribution of BLIPScores for ranked image-caption pairs. The box plot for rank 1 (higher-ranked captions) exhibits the highest median BLIPScore and a narrower score range, indicating better quality and consistency. In contrast, as the rank increases (i.e., captions are ranked lower), the median BLIPScore decreases, and the score range widens, reflecting lower quality and greater variability. Outliers suggest that some captions within certain ranks deviate significantly—either positively or negatively—from the majority. Overall, the figure reveals that BLIPScores decline with rank, highlighting the superior quality of higher-ranked captions. The similarity in medians for the first two ranks further supports focusing on these for tasks like fusion, as including lower-ranked captions may reduce overall accuracy.
Since the median BLIPScore value for the top three captions in the ranking is quite similar, we conduct an ablation study in which we combine the top three captions in the ranking instead of the top two. The model is kept in its original form except for the prompt, which is adapted to include three captions instead of two. The results reported in row 3 of Table 7 show how the use of three captions instead of two (whose results are in the last row) leads to a slight degradation in performance across most metrics. In particular, although the BLEU-4 and SPICE scores decrease only marginally, more noticeable drops are observed for METEOR, CIDEr, CAPTURE, and Polos. These results suggest that incorporating an additional caption in the prompt may introduce redundancy or conflicting semantic content that hinders the model’s ability to generate coherent and high-quality outputs. Overall, the configuration with two captions appears to provide a better balance between diversity and relevance, resulting in superior caption quality.
In Figure 10, we present the percentage of SoTA model pairs selected during the ranking phase. In practice, these results demonstrate which models most frequently belong to the captions that are fused by our model. As shown, the pair of models BLIP-2 and OFA is selected for 22% of the test image captions, while GIT and OFA are selected for 19%. The pair consisting of ExpNet-v2 and ViT-GPT2 accounts for only 2.4% of the generated captions.
Figure 9: Distribution of BLIPScore for image-caption pairs grouped by rank.
4.4.2 Caption fusion through LLM alternatives
To evaluate the generality and practicality of our fusion-based strategy, we replaced the proprietary Davinci model with two widely used open-weight LLMs: LLaMA-3.3-70B-Instruct and Mixtral-8x7B-Instruct-v0.1. These models were selected for their public availability and favorable performance-cost trade-offs. In all cases, we used the same “Sem-Syn Merge” prompt (see Section 4.4.3), and fused two captions per image for consistency across conditions.
The results, shown in rows 4 and 5 of Table 7, reveal significant performance differences between these alternatives. In particular, the variant using LLaMA shows a sharp decline across all quality metrics (B@4: 10.6, C: 16.3, S: 14.9), indicating that this model struggled to synthesize coherent and relevant captions using our fusion prompt. This degradation suggests that, despite its architectural size, LLaMA may lack robustness or instruction-following capabilities sufficient for the nuanced semantic alignment required in this task. Conversely, the Mixtral variant delivers much stronger results, achieving 12.7 in B@4, 27.1 in METEOR, and 23.4 in SPICE. While these scores are still lower than those obtained with Davinci, they demonstrate Mixtral’s ability to understand and recombine structured inputs meaningfully. Notably, Mixtral matches Davinci in CAPTURE (0.42) and approaches its performance in the Polos metric (0.70 vs. 0.74), suggesting that it can produce semantically grounded captions even with simpler or more cost-efficient architectures.
Beyond quality, we also consider inference time and cost-efficiency. According to benchmark data from Artificial Analysis [^1] https://artificialanalysis.ai/, Mixtral and LLaMA significantly outperform Davinci in latency, requiring only 6.6 and 5.5 seconds respectively, to generate 500 tokens, compared to the 18 seconds required by Davinci. This is particularly relevant for large-scale or real-time applications. In terms of monetary costs, the savings are substantial. Estimated total inference costs for the MS-COCO test set are \$18.50 for Davinci, versus just \$0.35 and \$0.30 for Mixtral and LLaMA, respectively—yielding over 98% cost reduction. These findings suggest that while Davinci remains the most effective for high-fidelity caption fusion, Mixtral offers a promising balance between performance and efficiency, and may be preferred in resource-constrained or deployment-oriented scenarios.
Table 7: Ablation experiments on the MS-COCO dataset.
#captions LLM Prompt B@4 M C S CAPTURE Polos
2 Davinci Basic 24.6 30.4 88.8 25.9 0.39 0.71
2 Davinci LLM-Optimized 29.3 31.2 102.1 24.9 0.42 0.73
3 Davinci Sem-Syn Merge 29.4 30.0 79.9 23.9 0.41 0.72
2 Llama Sem-Syn Merge 10.6 23.1 16.3 14.9 0.38 0.58
2 Mixtral Sem-Syn Merge 12.7 27.1 21.9 23.4 0.42 0.70
2 Davinci Sem-Syn Merge 30.0 31.5 106.0 26.5 0.42 0.74
4.4.3 Definition of the optimal prompt
  • Explicitly instructing the model to fuse sentences both syntactically and semantically facilitates the creation of a result that goes beyond a mere summary of the two captions. Instead, it produces a genuine amalgamation of information while eliminating redundancy in terms and/or content. For example, if both analyzed sentences feature the term “kite” in a specific syntactic role, the model comprehends that there is a single subject and avoids duplicating the term “kite” when generating the final caption.
Among all the attempts made, three representative prompts are reported below. The first is the simplest, manually engineered by the authors, aiming at a basic combination of the two input captions. The second is a more elaborated version, still manually designed by the authors, which includes explicit instructions to consider both semantic and syntactic aspects of the sentences. Finally, the third prompt was automatically generated by an LLM. To generate it, the LLM was provided with a meta-prompt that incorporated all the key aspects learned from the previous manual attempts – such as emphasizing the need for semantic and syntactic combination, resolving coreference across captions (e.g., “child” and “young girl”), and avoiding redundancy in the final output. The resulting LLM-generated prompt is more refined and explicitly guides the model to produce a cohesive and precise ensemble caption. The three prompts are reported below:
  • Basic – ‘Combine the meaning of these 2 sentences in 1 sentence. The sentences are: caption1; caption2. I want to get the best caption of the image”;
  • Semantic-Syntactic Merge – ‘Combine the meaning of these 2 sentences into 1 sentence, considering the semantic meaning and the syntactic meaning. The sentences are: caption1; caption2. These sentences describe an image, I want to get the best caption of the image, using the information in these two sentences”;
  • LLM-Optimized – ‘Analyze the following two sentences, both describing the same input image. Syntactically and semantically combine the information and meaning from both captions into a single, cohesive, and non-redundant description. Prioritize accuracy and conciseness, ensuring that shared entities are represented by the most precise term, and duplicated information is unified. The goal is a true ensemble caption that integrates all relevant details without repetition. Caption 1: caption1. Caption 2: caption2”.
Results for the model involving the Basic and LLM-Optimized prompts are reported in rows 1 and 2, while the results achieved by the final version involving the Sem-Syn Merge prompt are reported in row 3 of Table 7. A clear performance progression can be observed as the prompt design becomes increasingly structured and tailored to the task. Starting from the Basic prompt, which yields BLEU-4 of 24.6 and CIDEr of 88.8, we observe a substantial improvement with the LLM-Optimized prompt (BLEU-4 of 29.3 and CIDEr of 102.1), confirming the effectiveness of prompt refinement—even when automatically generated by another LLM. The best results across all evaluation metrics are obtained with the Sem-Syn Merge prompt, which was manually engineered to incorporate both semantic and syntactic alignment cues. This version achieves the highest scores in BLEU-4 (30.0), METEOR (31.5), CIDEr (106.0), SPICE (26.5), CAPTURE (0.42), and Polos (0.74). These findings demonstrate that prompt design has a significant impact on caption fusion quality, and that explicit instructions encouraging both semantic integration and syntactic fluency can effectively guide the model toward more accurate and informative captions. The results also highlight that, in this setting, careful human prompt engineering can outperform LLM-generated prompts.